A multistage reciprocating compressor runs at 900 rpm and delivers 3 kg per minute of air. It takes in air at 1 bar, 25° C and compresses it to 20 bar. All pistons have a stroke of 170 min.Calculate the index of compression?


1
Expert's answer
2021-06-28T08:54:11-0400

"\dfrac{T_2}{T_1}= (\dfrac{P_2}{P_1})^{\frac{n-1}{n}}"


"\dfrac{n-1}{n} = \dfrac{\ln{T_2/T_1}}{\ln{P_2/P_1}}"


"\dfrac{1}{n} = 1-\dfrac{\ln{25°C/25°C}}{\ln{20/1}}"


"\dfrac{1}{n} = 1-\dfrac{\ln1}{\ln20}"


"\dfrac1n= 1"


"\therefore n=1"